# The essential regularity of singular connections in Geometry

### v3

@@ 2. Statement of results
The proof of Theorem 2.5 is based on an asymmetry between the connection transformation law and the tensor transformation law for the curvature, noting that only the former involves derivatives of the Jacobian. To authors’ knowledge, Theorem 2.5 gives the first definitive theory for determining when a singularity, interpreted as a connection represented below its essential regularity, is removable or not. As far as we know, before this, there were no definitive conditions for determining whether or not connections with $L^{p}$ components could be regularized by coordinate transformation, nor was the essential regularity of a connection understood to be a geometric property of connections.
2.1. Discussion of results
Authors’ research program began with the problem of singular connections (below their essential regularity) which arose in the theory of GR shock waves [4, 14, 17, 26, 34, 37]. To construct local examples of connections with essential regularity $W^{m,p}$, simply take any collection of $W^{m,p}$ functions $\Gamma$ (viewed as the components of a matrix valued differential form) such that their exterior derivative is one derivative less regular than these functions. Conversely, it is straightforward to construct examples of connections below essential regularity by transforming the above $\Gamma$ according to the connection transformation law by a Jacobian of regularity below $W^{m,p}$.
@@ 2.1. Discussion of results
Theorem 2.3 (i) tells us that a connection defined by an atlas $\mathcal{A}_{s}$ globally exhibits one additional derivative of regularity in a transformed subatlas $\mathcal{A}_{s+1}$ which lies in $\mathcal{A}^{\text{max}}_{s}$, (the maximal $W^{2,p}$ extension of $\mathcal{A}_{s}$), if and only if ${\rm Riem}(\Gamma)$ has the same regularity as $\Gamma$ in $\mathcal{A}_{s}$, i.e., ${\rm Riem}(\Gamma)\in W^{s,p}_{\mathcal{A}_{s}}$. In fact, by applying the existence theory for the RT-equations, we construct the subatlas $\mathcal{A}_{s+1}$ within the smaller atlas $\mathcal{A}_{s}^{\text{max}(s+2)}\subset\mathcal{A}^{\text{max}}_{s}$, (see Section 6). The coordinate transformations between atlas $\mathcal{A}_{s}$ and $\mathcal{A}_{s+1}$ preserve the connection regularity $\Gamma\in W^{s,p}$, so by Lemma 2.1 the transition maps between coordinate systems in $\mathcal{A}_{s}$ and $\mathcal{A}_{s+1}$ have minimal regularity $W^{s+2,p}$, which implies $\mathcal{A}_{s+1}$ actually lies within the smaller atlas $\mathcal{A}_{s}^{\text{max}(s+2)}$. Theorem 2.3 also tells us that if ${\rm Riem}(\Gamma)$ is one derivative less regular than $\Gamma$ in $W^{s,p}$ in atlas $\mathcal{A}_{s}$, then no subatlas exists within the entire maximal $W^{2,p}$ extension $\mathcal{A}^{\text{max}}_{s}$ of $\mathcal{A}_{s}$ in which the connection is one derivative more regular than $W^{s,p}$.
@@
Theorem 2.3 (ii) determines an algorithm for lifting any connection to its essential regularity by use of the RT-equations, provided that $ess_{\mathcal{M}}(\Gamma)<\infty$. To explain this, start with a connection $\Gamma$ with regularity $W^{s_{0},p}$, defined in a given initial atlas $\mathcal{A}_{s_{0}}$, so $\Gamma\in W^{s_{0},p}_{\mathcal{A}_{s_{0}}}$, $s_{0}\geq 0,$ $p>n.$ By Theorem 2.3 (i), ${\rm Riem}(\Gamma)\in W^{s_{0},p}_{\mathcal{A}_{s_{0}}}$ implies the existence of atlas $\mathcal{A}_{s_{0}+1}$ in the maximal $W^{2,p}$ extension $\mathcal{A}^{\text{max}}_{s_{0}}$ of $\mathcal{A}_{s_{0}}$, [^7] Note, as discussed below Theorem 2.3, we actually have $\mathcal{A}_{s_{0}+1}\subset\mathcal{A}^{\text{max}(s_{0}+2,p)}_{s_{0}}\subset\mathcal{A}^{\text{max}}_{s_{0}}$, but to keep the discussion simple we consider here $\mathcal{A}^{\text{max}}$. constructed by solving the RT-equations, such that $\Gamma\in W^{s_{0}+1,p}_{\mathcal{A}_{s_{0}+1}}$. Letting $\Gamma_{s}$ and ${\rm Riem}(\Gamma_{s})$ denote the components of $\Gamma$ and ${\rm Riem}(\Gamma)$ in atlas $\mathcal{A}_{s}$, respectively, and applying Theorem 2.3 to $\Gamma_{s_{0}+1},$ we see that $\Gamma_{s_{0}+1}$ can be further regularized by coordinate transformation if and only if the components ${\rm Riem}(\Gamma_{s_{0}+1})$ have the regularity of components $\Gamma_{s_{0}+1}$, i.e., if and only if ${\rm Riem}(\Gamma)\in W^{s_{0}+1,p}_{\mathcal{A}_{s_{0}+1}}$. If so, Theorem 2.3 (i) asserts that a new subatlas $\mathcal{A}_{s_{0}+2}$ of $\mathcal{A}^{\text{max}}_{s_{0}}$ can be constructed from $\mathcal{A}_{s_{0}+1}$ by solving the RT-equations, such that $\Gamma\in W^{s_{0}+2,p}_{\mathcal{A}_{s_{0}+2}}$. It follows, then, that this procedure continues until one reaches the maximal value $s=m\equiv ess_{\mathcal{M}}(\Gamma)\geq s_{0}$, which by Theorem 2.3 must be the first level of regularity $W^{m,p}$, $m\geq s_{0}$, at which ${\rm Riem}(\Gamma_{m})$ is one full derivative less regular than $\Gamma_{m}$ in atlas $\mathcal{A}_{m}$. Theorem 2.5 asserts that this value of $m$ will always be the essential regularity of $\Gamma$ as defined in Definition 2.2; and the composition of coordinate transformations determined by the RT-equations at each step $s=s_{0},...,m$, provides coordinate transformations which lift the connection $\Gamma$ defined in the original atlas $\mathcal{A}_{s_{0}}$, directly to the atlas $\mathcal{A}_{m}$ in which $\Gamma$ achieves its essential regularity. In this case, for each $s=s_{0},...,m,$ the atlas $\mathcal{A}_{s}$ is a subatlas of the maximal $W^{2,p}$ extension $\mathcal{A}^{\text{max}}_{s_{0}}$ of the original atlas $\mathcal{A}_{s_{0}}$, and the regularity of atlas $\mathcal{A}_{s}$ will be $W^{s+2,p}$, the lowest regularity required to preserve the $W^{s,p}$ regularity of $\Gamma_{s}$ under coordinate transformations within the atlas $\mathcal{A}_{s}$. In the case when $ess_{\mathcal{M}}(\Gamma)=m=\infty$, this procedure will regularize the connection components to $W^{s,p}$ for every integer $s<\infty$, but, since author’s existence for the RT-equations requires restrictions to smaller domains, not necessarily all the way to $s=\infty$.
In Section 7 we establish local and global estimates on connections under a one- and two-step regularization. Sobolev regularity of a tensor or connection is invariant under sufficiently regular coordinate transformations, but the value of the Sobolev norm itself is not invariant except in the case of a Riemannian metric, making the interpretation of Sobolev estimates problematic. Our theory provides an invariant meaning to the essential Sobolev regularity of a connection, but again the value of the Sobolev norm still depends on the coordinate system. Note however, the local regularization of singular connections by the RT-equations does imply $W^{s+1,p}$ estimates on regularized connections together with the regularizing coordinate transformations, in terms of the $W^{s,p}$-norms of connection and curvature in starting coordinates. But the RT-equations provide no bound on the regularization of the curvature in Sobolev norms above $W^{s,p}$. So we do not get any additional estimates in terms of the original connection and curvature under a regularization by the RT-equations requiring more than two steps to essential regularity.
@@ 2.2. Hierarchy of $C^{\infty}$ differentiable structures
Authors find it interesting that naively one might think that the regularity of a manifold is given by the regularity of its atlas, but this cannot be a geometric property of manifolds alone, because atlases can be extended to a unique maximal $C^{1}$ atlas, and all subsequent restrictions to $C^{\infty}$ subatlases are $C^{\infty}$ equivalent in the sense of [16, Thm. 2.9]. Thus a manifold alone does not have enough structure to determine a geometric level of regularity. Our results here establish that a manifold together with a connection does have an inherent geometric level of regularity, the essential regularity of the connection.
@@ 3. The RT-equations
In this section we establish a refined existence theory for the RT-equations which is required for the results in this paper. We begin by recording authors’ prior results on the Regularity Transformation (RT-) equations, developed in a series of papers [28, 29, 30, 31, 32, 33, 27], as required for this paper. [^8] The RT-equations are referred to as the Reintjes-Temple equations in [30]. We proved in [29, 31] that solutions of the RT-equations furnish local coordinate transformations which regularize connections to one derivative of regularity above their Riemann curvature, as measured in a starting coordinate system. In prior papers we were content with the one step regularization of connections to one derivative above the initial curvature regularity, and considered the generic case when the curvature regularity stays fixed under the regularizing transformation. From the point of view of the present paper, we did not take account of the fact that the curvature is regularized along with the connection when connection regularity is more than one derivative below essential regularity, as asserted by Theorem 2.3 above. In our prior papers we did not fully understand that there can be an automatic regularization of the curvature and that subsequent use of the RT-equations would take one to an essential, i.e., best possible, geometric regularity of a connection. [^9] In this paper we thus distinguish “essential regularity” from “optimal regularity”, the latter being standard terminology (cf. [18, 3]) we used in our prior papers. That is, we said a connection has been lifted to “optimal regularity” by coordinate transformation if it exhibits one derivative of regularity above its Riemann curvature in some starting coordinate system–a coordinate dependent terminology which is imprecise in light of our new results here.
@@ 3.1. Derivation of the RT-equations
where $\Delta\equiv\delta d+d\delta=\partial_{x^{0}}^{2}+...+\partial_{x^{n}}^{2}$ is the Euclidean Laplacian, $\langle\cdot\;$$\cdot\rangle$ is a matrix-valued inner product and $\wedge$ the wedge product on matrix valued differential forms, (see [28, Ch.3] or [31, Ch.5] for detailed definitions). At this stage, equations (3.2) - (3.3) neither appear solvable, nor need a solution $J$ be a true Jacobian that is integrable to coordinates, i.e. satisfying ${\rm Curl}(J)=0$. To complete the equations, view $A\equiv\delta\tilde{\Gamma}$ as a free matrix valued function—this choice was motivated by the Riemann-flat condition for optimal regularity in [27], because the latter only involves $d\tilde{\Gamma}$, but not $\delta\tilde{\Gamma}$. Substituting $A\equiv\delta\tilde{\Gamma}$ in (3.2) - (3.3), and viewing $A$ as a new unknown matrix valued function, we next impose on equation (3.3) the condition ${\rm Curl}(J)=0$ for integrability. For this, we introduce the vectorization $\vec{J}^{\mu}=J^{\mu}_{\nu}dx^{\nu}$ of $J$, so that ${\rm Curl}(J)\equiv d\vec{J}$, and impose equivalently $d\vec{J}=0$ as the integrability condition. By a fortuitous cancellation the regularities in different terms of the equation become consistent, and the computations in [28] eventually lead to the RT-equations:
@@ 3.2. How the RT-equations yield optimal regularity
Our iteration scheme, which is based on solving the linear Poisson equation at each stage, applies to the reduced RT-equations (3.8) - (3.10) at the low regularity of $L^{p}$ connections with $d\Gamma\in L^{p/2}$, and locally establishes existence of solutions $(J,B)$, (i.e, in neighborhoods $\Omega^{\prime}$ of points), such that $J$ is point-wise an invertible matrix, cf. [31, Thm 6.4].
That any solution $J$ is a Jacobian integrable to coordinates is a built-in property of (3.8) - (3.10), provided that the integrability condition $d\vec{J}\equiv{\rm Curl}(J)=0$ holds on the boundary $\partial\Omega^{\prime}$, c.f. [31, Thm 6.4]. That is, combining (3.8) with (3.9), a computation shows that $\omega\equiv d\vec{J}$ is a solution of Laplace’s equation $\Delta\omega=0$, which together with our boundary data implies that $\omega=0$ throughout $\Omega^{\prime}$. This implies that $J$ is integrable to coordinates.
@@ 4. The Link Between Connection and Jacobian Regularity
Let $J$ be the Jacobian of a coordinate transformation which transforms $\Gamma_{y}\in W^{s,p}$ to $\Gamma_{x}$, and let $0\leq r\leq s$. Then the components of $\Gamma_{x}$ satisfy $\Gamma_{x}\in W^{r,p}$ if and only if the components of $J$ satisfy $J\in W^{r+1,p}$.
@@ 5. The Local Theory of Essential Regularity
$(\Gamma_{x})^{\mu}_{\rho\nu}=(J^{-1})^{\mu}_{\alpha}\Big(J^{\beta}_{\rho}J^{\gamma}_{\nu}\,(\Gamma_{y})^{\alpha}_{\beta\gamma}+\tfrac{\partial}{\partial x^{\rho}}J^{\alpha}_{\nu}\Big).$
@@ 7. Estimates for Two-Step Regularizations
The one-step regularization by the RT-equations provides estimates on the regularized connections and the regularizing coordinate transformations in terms of the $W^{s,p}$-norm of the curvature in the starting coordinates, cf. Theorems 3.1 and 3.2. This can be extended to estimates of a two-step regularization when the curvature in the starting coordinates is one derivative more regular than the connection, (cf. Case (3) of Theorem 2.5), by using the corresponding higher $W^{s+1,p}$-norm on the curvature in the starting coordinates, because the curvature regularity stays fixed under a one- or two-step regularization. However, the RT-equations provide no estimates for the regularization of the curvature, and thus do not provide estimates for the connection when lifted to essential regularity by more than two steps (which requires applying the RT-equations to the implicitly regularized curvature). The estimates for the two-step regularization in the local case is stated in Corollary 7.1 below. This is extended to a global setting on compact manifolds in Corollaries 7.2 and 7.3 below. We assume throughout that $p>n\geq 2$, $s\geq 0$.
@@
The estimates of Theorems 3.1 and 3.2 extend to global estimates, provided the incoming bounds (7.1) hold in a suitable uniform sense over suitably uniform domains. We implement this uniformity by assuming $\mathcal{M}$ to be a compact manifold. For one-step regularizations we have the following result.
@@
Assuem $\Gamma\in W^{s,p}_{\mathcal{A}_{s}}$ for some $W^{s+2,p}$ atlas $\mathcal{A}_{s}$ on a compact manifold $\mathcal{M}$ such that
@@
This research was supported by CityU Start-up Grant (7200748), CityU Strategic Research Grant (7005839) and General Research Fund (ECS 21306524).

### v4

@@ 2. Statement of results
The proof of Theorem 2.5 is based on an asymmetry between the connection transformation law and the tensor transformation law for the curvature, noting that only the former involves derivatives of the Jacobian. To the authors’ knowledge, Theorem 2.5 gives the first definitive theory for determining when a singularity, interpreted as a connection represented below its essential regularity, is removable or not. As far as we know, before this, there were no definitive conditions for determining whether or not connections with $L^{p}$ components could be regularized by coordinate transformation, nor was the essential regularity of a connection understood to be a geometric property of connections.
2.1. Discussion of Theorem 2.3
@@
Part (i) of Theorem 2.3 asserts that a connection defined by an atlas $\mathcal{A}_{s}$ globally exhibits one additional derivative of regularity in a transformed subatlas $\mathcal{A}_{s+1}$ which lies in $\mathcal{A}^{\text{max}}_{s}$, (the maximal $W^{2,p}$ extension of $\mathcal{A}_{s}$), if and only if ${\rm Riem}(\Gamma)$ has the same regularity as $\Gamma$ in $\mathcal{A}_{s}$, i.e., ${\rm Riem}(\Gamma)\in W^{s,p}_{\mathcal{A}_{s}}$. Conversely, Theorem 2.3 asserts that if ${\rm Riem}(\Gamma)$ is one derivative less regular than $\Gamma$ in $W^{s,p}$ in atlas $\mathcal{A}_{s}$, then no subatlas exists within the entire maximal $W^{2,p}$ extension $\mathcal{A}^{\text{max}}_{s}$ of $\mathcal{A}_{s}$ in which the connection is one derivative more regular than $W^{s,p}$.
@@
Part (ii) of Theorem 2.3 determines an algorithm for lifting any connection to its essential regularity by use of the RT-equations, provided that $ess_{\mathcal{M}}(\Gamma)<\infty$. To explain this, start with a connection $\Gamma$ with regularity $W^{s_{0},p}$ below its essential regularity, defined in a given initial atlas $\mathcal{A}_{s_{0}}$, so $\Gamma\in W^{s_{0},p}_{\mathcal{A}_{s_{0}}}$, $s_{0}\geq 0,$ $p>n.$ By Theorem 2.3 (i), ${\rm Riem}(\Gamma)\in W^{s_{0},p}_{\mathcal{A}_{s_{0}}}$ implies the existence of atlas $\mathcal{A}_{s_{0}+1}$ in the maximal $W^{2,p}$ extension $\mathcal{A}^{\text{max}}_{s_{0}}$ of $\mathcal{A}_{s_{0}}$, [^8] As discussed in Footnote 7, we actually have $\mathcal{A}_{s_{0}+1}\subset\mathcal{A}^{\text{max}(s_{0}+2,p)}_{s_{0}}\subset\mathcal{A}^{\text{max}}_{s_{0}}$, but to keep the discussion simple we consider here $\mathcal{A}^{\text{max}}$. constructed by solving the RT-equations, such that $\Gamma\in W^{s_{0}+1,p}_{\mathcal{A}_{s_{0}+1}}$. Letting $\Gamma_{s}$ and ${\rm Riem}(\Gamma_{s})$ denote the components of $\Gamma$ and ${\rm Riem}(\Gamma)$ in atlas $\mathcal{A}_{s}$, respectively, and applying Theorem 2.3 to $\Gamma_{s_{0}+1},$ we see that $\Gamma_{s_{0}+1}$ can be further regularized by coordinate transformation if and only if the components ${\rm Riem}(\Gamma_{s_{0}+1})$ have the regularity of components $\Gamma_{s_{0}+1}$, i.e., if and only if ${\rm Riem}(\Gamma_{s_{0}+1})\in W^{s_{0}+1,p}_{\mathcal{A}_{s_{0}+1}}$. If so, Theorem 2.3 (i) asserts that a new subatlas $\mathcal{A}_{s_{0}+2}$ of $\mathcal{A}^{\text{max}}_{s_{0}}$ can be constructed from $\mathcal{A}_{s_{0}+1}$ by solving the RT-equations, such that $\Gamma\in W^{s_{0}+2,p}_{\mathcal{A}_{s_{0}+2}}$. It follows, then, that this procedure continues until one reaches the maximal value $s=m\equiv ess_{\mathcal{M}}(\Gamma)\geq s_{0}$, which by Theorem 2.3 must be the first level of regularity $W^{m,p}$, $m\geq s_{0}$, at which ${\rm Riem}(\Gamma_{m})$ is one full derivative less regular than $\Gamma_{m}$ in atlas $\mathcal{A}_{m}$.
Theorem 2.5 asserts that this value of $m$ will always be the essential regularity of $\Gamma$ as defined in Definition 2.2; and the composition of coordinate transformations determined by the RT-equations at each step $s=s_{0},...,m$, provides coordinate transformations which lift the connection $\Gamma$ defined in the original atlas $\mathcal{A}_{s_{0}}$, directly to the atlas $\mathcal{A}_{m}$ in which $\Gamma$ achieves its essential regularity. In this case, for each $s=s_{0},...,m,$ the atlas $\mathcal{A}_{s}$ is a subatlas of the maximal $W^{2,p}$ extension $\mathcal{A}^{\text{max}}_{s_{0}}$ of the original atlas $\mathcal{A}_{s_{0}}$, and the regularity of atlas $\mathcal{A}_{s}$ will be $W^{s+2,p}$, the lowest regularity required to preserve the $W^{s,p}$ regularity of $\Gamma_{s}$ under coordinate transformations within the atlas $\mathcal{A}_{s}$. In the case when $ess_{\mathcal{M}}(\Gamma)=m=\infty$, this procedure will regularize the connection components to $W^{s,p}$ for every integer $s<\infty$, but, since the existence theory for the RT-equations requires restrictions to smaller domains, not necessarily all the way to $s=\infty$.
@@ 2.2. Hierarchy of $C^{\infty}$ differentiable structures
The authors find it interesting that naively one might think that the regularity of a manifold is given by the regularity of its atlas, but this cannot be a geometric property of manifolds alone, because atlases can be extended to a unique maximal $C^{1}$ atlas, and all subsequent restrictions to $C^{\infty}$ subatlases are $C^{\infty}$ equivalent in the sense of [16, Thm. 2.9]. Thus a manifold alone does not have enough structure to determine a geometric level of regularity. Our results here establish that a manifold together with a connection does have an inherent geometric level of regularity, the essential regularity of the connection.
@@ 3. The RT-equations
In this section we establish a refined existence theory for the RT-equations which is required for the results in this paper. We begin by recording the authors’ prior results on the Regularity Transformation (RT-) equations, developed in a series of papers [29, 30, 31, 32, 33, 34, 28], as required for this paper. We proved in [30, 32] that solutions of the RT-equations furnish local coordinate transformations which regularize connections to one derivative of regularity above their Riemann curvature, as measured in a starting coordinate system. In prior papers we were content with the one step regularization of connections to one derivative above the initial curvature regularity, and considered the generic case when the curvature regularity stays fixed under the regularizing transformation. From the point of view of the present paper, we did not take account of the fact that the curvature is regularized along with the connection when connection regularity is more than one derivative below essential regularity, as asserted by Theorem 2.3 above. In our prior papers we did not fully understand that there can be an automatic regularization of the curvature and that subsequent use of the RT-equations would take one to an essential, i.e., best possible, geometric regularity of a connection. [^9] In this paper we thus distinguish “essential regularity” from “optimal regularity”, the latter being standard terminology (cf. [18, 3]) we used in our prior papers. That is, we said a connection has been lifted to “optimal regularity” by coordinate transformation if it exhibits one derivative of regularity above its Riemann curvature in some starting coordinate system–a coordinate dependent terminology which is imprecise in light of our new results here.
@@ 3.1. Derivation of the RT-equations
where $\Delta\equiv\delta d+d\delta=\partial_{x^{0}}^{2}+...+\partial_{x^{n}}^{2}$ is the Euclidean Laplacian, $\langle\cdot\;$$\cdot\rangle$ is a matrix-valued inner product and $\wedge$ the wedge product on matrix valued differential forms, (see [29, Ch.3] or [32, Ch.5] for detailed definitions). At this stage, equations (3.2) - (3.3) neither appear solvable, nor need a solution $J$ be a true Jacobian that is integrable to coordinates, i.e. satisfying ${\rm Curl}(J)=0$. To complete the equations, view $A\equiv\delta\tilde{\Gamma}$ as a free matrix valued function—this choice was motivated by the Riemann-flat condition for optimal regularity in [28], because the latter only involves $d\tilde{\Gamma}$, but not $\delta\tilde{\Gamma}$. Substituting $A\equiv\delta\tilde{\Gamma}$ in (3.2) - (3.3), and viewing $A$ as a new unknown matrix valued function, we next impose on equation (3.3) the condition ${\rm Curl}(J)=0$ for integrability. For this, we introduce the vectorization $\vec{J}^{\mu}=J^{\mu}_{\nu}dx^{\nu}$ of $J$, so that ${\rm Curl}(J)\equiv d\vec{J}$, and impose equivalently $d\vec{J}=0$ as the integrability condition. By a fortuitous cancellation the regularities in different terms of the equation become consistent, and the computations in [29] eventually lead to the RT-equations:
@@ 3.2. How the RT-equations yield optimal regularity
Our iteration scheme, which is based on solving the linear Poisson equation at each stage, applies to the reduced RT-equations (3.8) - (3.10) at the low regularity of $L^{p}$ connections with $d\Gamma\in L^{p/2}$, and locally establishes existence of solutions $(J,B)$, (i.e, in neighborhoods $\Omega^{\prime}$ of points), such that $J$ is point-wise an invertible matrix, cf. [32, Thm 6.4].
That any solution $J$ is a Jacobian integrable to coordinates is a built-in property of (3.8) - (3.10), provided that the integrability condition $d\vec{J}\equiv{\rm Curl}(J)=0$ holds on the boundary $\partial\Omega^{\prime}$, c.f. [32, Thm 6.4]. That is, combining (3.8) with (3.9), a computation shows that $\omega\equiv d\vec{J}$ is a solution of Laplace’s equation $\Delta\omega=0$, which together with our boundary data implies that $\omega=0$ throughout $\Omega^{\prime}$. This implies that $J$ is integrable to coordinates.
@@ 4. The Link Between Connection and Jacobian Regularity
Let $J$ be the Jacobian of a coordinate transformation which transforms $\Gamma_{y}\in W^{s,p}$ to $\Gamma_{x}$, and let $0\leq r\leq s$ and $p>n$. Then the components of $\Gamma_{x}$ satisfy $\Gamma_{x}\in W^{r,p}$ if and only if the components of $J$ satisfy $J\in W^{r+1,p}$.
@@ 5. The Local Theory of Essential Regularity
Based on Theorem 5.2 it is now easy to understand how to produce a connection below its essential regularity. This is based on the following asymmetry between the connection transformation law (5.2) and the tensor transformation law (5.1): (5.2) involves Jacobian derivatives, while (5.1) involves only the undifferentiated Jacobian. It then follows directly by Lemma 4.1 that transforming a $W^{m,p}$ connection $\Gamma$ by a Jacobian of regularity below $W^{m,p}$ always produces a connection below its essential regularity. Conversely, to construct examples of connections with essential regularity $W^{m,p}$ locally, simply take any collection of $W^{m,p}$ functions $\Gamma$, (viewed as the components of a matrix valued differential form), such that their exterior derivative (and hence their Riemann curvature) is one derivative less regular than these functions.
@@ 7. Estimates for Two-Step Regularizations
In this section, we establish local and global estimates on Sobolev norms of connections under a one- and two-step regularization. The one-step regularization by the RT-equations provides estimates on the regularized connections and the regularizing coordinate transformations in terms of the $W^{s,p}$-norm of the curvature in the starting coordinates, cf. Theorems 3.1 and 3.2. This can be extended to estimates of a two-step regularization when the curvature in the starting coordinates is one derivative more regular than the connection, (cf. Case (3) of Theorem 2.5), by using the corresponding higher $W^{s+1,p}$-norm on the curvature in the starting coordinates, because the curvature regularity stays fixed under a one- or two-step regularization. However, the RT-equations provide no estimates for the regularization of the curvature. Thus they do not provide estimates for the connection when lifted to essential regularity by more than two steps, because this would require applying the RT-equations to the implicitly regularized curvature. Hence estimates for multi-step regularizations (more than two) appear out of reach.
The estimates for the two-step regularization in the local case is stated in Corollary 7.1 below. This is extended to a global setting on compact manifolds in Corollaries 7.2 and 7.3 below. We assume throughout that $p>n\geq 2$, $s\geq 0$.
@@
Before extending Corollary 7.1 to a global setting, note that Sobolev regularity of a tensor or connection is invariant under sufficiently regular coordinate transformations, but the value of the Sobolev norm itself is not invariant, except in the special case of a Riemannian metric. We address this problem here by basing norms on the Euclidean metric in each coordinate chart over a compact manifold. That is, Corollary 7.1 extends to global estimates, provided the incoming bound holds in a suitable uniform sense over suitably uniform domains, which we implement by assuming $\mathcal{M}$ to be a compact manifold. For one-step regularizations we have the following result.
@@
M.R. was supported by the City University of Hong Kong (Start-up Grant 7200748 and CityU Strategic Grant 7005839) and by the Hong Kong Research Grants Council (ECS-Grant 21306524 and GRF-Grant 11303326).